In ohms law what is slope of I-V and V-I graph represents? both represents resistance or one represents capacitance and one represents resistance?

Grade:12th pass

1 Answers

Arun
6 years ago
From Ohm's law- V= IR
where, V = Voltage, I = current , R = Resistance
So, Slope of I-V graph = V/I = R 
and the slope of V-I graph = I/V = 1/R = Conductance.